A L^{A}TEX Package to Prepare Author and Affiliation Blocks

Patrick W. Daly

This paper describes L^{A}TEX 2ε package authblk version 1.3 from 2001/02/27 It is part of the preprint collection of packages

Summary

The stripped version of this file contains the following brief description:

% A LaTeX2e package to redefine the \author command to work as normal or
% to allow a footnote style of author/affiliation input.
%
% Three methods of input:
%  1. Standard LaTeX
%     \author{Name1//Affil1 /and Name2 and Name3//Affil2 . . .}
% (use \authorcr, not \\, to force new line within the author text)
%  2. Automatic mode (selects modes 1 or 3 depending on number of affiliations)
%     \author{Name1}
%     \affil{Affil1}
%     \author{Name2}
%     \author{Name3}
%     \affil{Affil2}
%     . . .
%  3. Footnotes explicitly
%     \author[*]{Name1}
%     \author[**]{Name2}
%     \author[**]{Name3}
%     \affil[*]{Affil1}
%     \affil[**]{Affil2}
%     . . .

1 Introduction

The standard L^{A}TEX syntax for entering author names and their affiliations is rather limited. The standard is to have author blocks consisting of all authors from one institute grouped together. However, if there are many institutes, or if the authors are ordered in some other way (e.g. alphabetically) it is more appropriate to use a footnote system. Both methods should be available for any paper since the style depends more on the number of authors and affiliations and their mixture.

This package offers a syntax for entering author names and their affiliations that permits both styles, and is also compatible with the standard L^{A}TEX syntax. It places its results in the control sequence \@author, which is the standard container for the author/affiliation names; in this way it can be used immediately with standard L^{A}TEX classes, or with packages that modify the title/author block.

2 Invoking the Package

The macros in this package are included in the main document with the \usepackage command of L^{A}TEX 2ε,

     \documentclass[..]{...}
     \usepackage[options]{authblk}

where the possible options are described below in Section 5.

3 Usage

Authors and affiliations can be printed in author blocks, as

J. B. Smith Institute for Research Hamilton, Ontario

C. T. Appleton and P. R. Bailey NASA, Greenbelt, MD

or in a line with footnote markers for the affiliations, as

J. B. Smith^{1}, C. T. Appleton^{2}, and P. R. Bailey^{2}

1Institute for Research, Hamilton, Ontario 2NASA, Greenbelt, MD

\author The user can enter the author and affiliation information in a number of dif- \affil ferent manners.

1. Standard L^{A}TEX, with syntax

         \author{. . . \and. . . \and. . . }

where each block of author(s)/affiliation comes between the \and commands, containing possibly \\ for new lines;

2. Automatic switching between author block or footnote style, information input with a series of \author, \affil commands as

         \author{author1 }
         \author{author2 }
         \affil{affil1 }
         \author{author3 }
         \author{author4 }
         \affil{affil2 }

⋯ where the authors will appear all in one (possibly continued) line with automatic footnotes, and the affiliations appear in separate lines below;

3. Footnote style with manual footnote markers given as optional arguments; in this case, the \author and \affil commands can be fully separated:

         \author[1]{author1 }
         \author[1]{author2 }
         \author[2]{author3 }
         \author[2]{author4 }
         \affil[1]{affil1 }
         \affil[2]{affil2 }

⋯

Method 1 can only produce author blocks; method 3 only footnotes, while 2 can produce both. If the number of affiliations is greater than Maxaffil, footnotes are used, otherwise author blocks.

\authorcr With method 1, the \\ command is used to switch to affiliation text, possibly changing the font; if the author text is to be forced to a new line, use \authorcr instead. Warning: methods 2 and 3 can only be used after authblk has been loaded, since \affil is otherwise not defined. However, a single \author command (method 1) can appear before \usepackage, or even before \documentclass, as is allowed in standard L^{A}TEX.

4 Configuration

Certain aspects of the author/affiliation blocks can be controlled by the user through redefinitions.

Maxaffil As already mentioned, the maximum number of affiliations before automatically switching from author block to footnote mode is set in the counter Maxaffil. Redefine this with

     \setcounter{Maxaffil}{2}

The default is 2.

\Authfont The author and affiliation lines can be set in different fonts, determined by these \Affilfont macros. By default they are \normalfont. Redefine them with \renewcommand,

for example:

      \renewcommand\Authfont{\scshape}
      \renewcommand\Affilfont{\itshape\small}

\affilsep The interline spacing between the author and affiliation lines is set by the length \affilsep. Change this with \setlength, as

     \setlength{\affilsep}{1em}

This spacing applies to both footnote and author block modes.

\Authsep In footnote mode, one can change the texts that comes between the authors’

\Authand
\Authands

names, between two authors, and before the last author when there are more than two. For example,

      \renewcommand\Authsep{, }
      \renewcommand\Authand{ and }
      \renewcommand\Authands{, and }

This example actually shows the defaults.

5 Options during Loading

Some reconfigurations are provided for by means of options to the \usepackage command. These are: blocks sets Maxaffil to ∞ so that the automatic selection will always be author/

affiliation blocks; noblocks sets Maxaffil to 0 so that the automatic selection will always be foot-

note mode; maxn (n = 2, 6) sets Maxaffil to n; auth-sc sets author font to small caps; auth-sc-lg sets author font to small caps, but with size \large; auth-lg sets author font to normal, but in size \large; affil-sl sets affiliation font to slanted; affil-it sets affiliation font to italic; german changes ‘and’ to ‘und’ in author lists.